OpenStudy (friendzone):

Write the equation of the line that passes through (2,-7) and is perpendicular to 5x+2y=18 in standard form.

OpenStudy (anonymous):

5x + 2y = 7 2y = -5x + 7 (sub. 5x from both sides) y = (-5/2)x + 7/2 (divide both sides by 2) Right here is where you use the slope to find the line parallel. Remember that if two lines are parallel, their slopes must be equal. This means any equation modeled by y = (-5/2)x + b will be parallel to the original, where b is the y-intercept. To find the line that passes through the point (-2,-4), substitute the respective x and y values into the general equation. This gives (-4) = (-5/2)(-2) + b -4 = 5 + b -9 = b Use this to find the specific equation of the line. Substitute b back in... y = (-5/2)x - 9 That is the correct line, but not an answer you gave. In order to find which equation is the correct answer, multiply both sides by 2 to get 2y = -5x - 18 Now add 5x to both sides and voila! 5x + 2y = -18

OpenStudy (anonymous):

5x + 2y = -18